OpenStudy (anonymous):

What is d over 3d to the 3rd power simplified (fraction)

OpenStudy (anonymous):

Do you mean: \(\dfrac{d}{3d^3}\) or \(\left(\frac{d}{3d}\right)^3\)

OpenStudy (anonymous):

The first one

OpenStudy (anonymous):

Is only the d cubed or the quantity 3d cubed?

OpenStudy (anonymous):

The quantity of d is cubed

OpenStudy (anonymous):

\(\dfrac{x^5}{x^2} = x^{5-2} = x^3\) When dividing the exponents, subtract the exponents.

OpenStudy (anonymous):

If you expand it out, it might make more sense.

OpenStudy (anonymous):

\(\dfrac{d}{3ddd}\) Reduce the fraction by dividing the numerator and the denominator by the same thing.

OpenStudy (anonymous):

Ok. Thank you